OpenStudy (anonymous):

solve the substitution method 5x-9y=-15 3x+31=y

OpenStudy (anonymous):

Substitute the second eqn (already equal to y) into the first: 5x - 9 * (3x + 31) = -15 5x - 27x - 279 = -15 -22x = 264 x = -12 Now, substitute x back into second equation: 3 * (-12) + 31 = y -36 + 31 = y y = -5

OpenStudy (anonymous):

so how do i get my answer?

OpenStudy (anonymous):

I would assume the answer needed was the x and y values: x=-12 and y=-5. You asked to solve by substitution, which the work shows.

OpenStudy (anonymous):

so how would i type it as a ordered pair?

OpenStudy (anonymous):

An ordered pair is just (x, y), so this problem would be (-12, -5).